OpenStudy (anonymous):

When 1.0 g of solid NaOH (Triangle Hsoln = –445.1 kJ/mol) dissolves in 10 L of water, how much heat is released? would it be 445.1?

OpenStudy (jfraser):

the 445kJ is per MOLE of NaOH, you're dissolving only 1 GRAM of NaOH...

OpenStudy (anonymous):

so then would it be.....11.1kJ?

OpenStudy (jfraser):

\[\frac{445kJ}{1mol NaOH}*\frac{1mol NaOH}{40g NaOH} = 11.1\frac{kJ}{g}\] yes

OpenStudy (anonymous):

thank you!!

OpenStudy (jfraser):

yw